a.   You are given the following information about a hypothetical economy;

C = 800 + 0.75Y

I = 500

G = 900

Where

C = Consumption

I = Investment

G = Government spending

(i) Compute the equilibrium level of national output and consumption.        

(ii) What is the size of government expenditure multiplier (1mk)

(iii) Interpret the government expenditure multiplier in (ii) above                

(iv) Differentiate between accelerator and multiplier effects.                       

(v) Explain the limitations to the application of multiplier in economic management.       


1
Expert's answer
2021-06-23T17:44:21-0400

(i) The equilibrium level of national output and consumption are:

Y = C + I + G = 800 + 0.75Y + 500 + 900,

0.25Y = 2,200,

Y = 8,800.

C = 800 + 0.75×8,800 = 7,400.

(ii) The size of government expenditure multiplier is:

g = G/Y = 900/8,800 = 0.102.

(iii) The government expenditure multiplier means that the increase in government spending will cause almost 10 times higher increase in GDP.

(iv) The multiplier describes the relationship between investment and income, i.e., the effect of investment on income. But in this concept, we are not concerned about the effect of income on investment. This effect is covered by the 'accelerator'.

(v) In actual practice, the operation of multiplier is affected by a number of factors. Given the MPC, the whole of the increment of income in each period may not be spent on consumption.